# Accurately simulating core-collapse self-interacting dark matter halos

### v2

@@ 3.7 Evolution of projected enclosed mass
Figure 9: Time evolution of the enclosed mass for our highest-resolved simulation (simulation Q in Tab. 1). The relative change of the enclosed mass within various radii is shown as a function of time. The enclosed mass is computed in projection, i.e. it is the mass within a cylinder of radius $r_{\mathrm{2D}}$.
@@
Now, we want to get closer to observations, in particular gravitational lensing, and measure the enclosed mass in projection. In Fig. 9, the enclosed mass within several radii is shown as a function of time for our highest-resolved simulation ($N=5\times 10^{7}$). For small radii, we find the enclosed mass to first decrease during the core expansion phase and later to increase when the gravothermal collapse dominates, as we also found in our three-dimensional analysis. This is in contrast to the bound mass of the system. The self-interactions create a few particles which have velocities exceeding the escape velocity, leading to a decreasing mass of the halo. Even more particles may gain enough kinetic energy to travel to fairly large radii, but still remain bound. In line with this, we find that for the largest radius shown in Fig. 9, the mass slightly decreases. Related to this, the velocity anisotropy at large radii increases, i.e. radial motion becomes more important, as we show in Appendix G.
@@
If one defines a radius $r_{\mathrm{eq}}$ as the one at which mass inflow and mass outflow have the same size, i.e. $\dot{M}(<r_{\mathrm{eq}})=0$. Then $r_{\mathrm{eq}}$ decreases with time when the halo is undergoing the gravothermal collapse. This implies that we do not expect the mass within a fixed radius, e.g. 1 kpc, to keep continuously increasing. Instead, it starts to decrease as we can see for a few of the radii that we have chosen for Fig. 9. The smaller the radius the later the point in time where the enclosed mass starts to decrease. [^1] For a more complicated model, such as the one studied by Patil & Fischer (2025), the decrease in mass at late stages can be avoided by approaching a stable solution with an increased enclosed mass compared to collisionless DM instead of undergoing a gravothermal catastrophe.
@@
Figure 10: Projected logarithmic density profile slope as a function of projected enclosed mass. For our highest-resolved simulation evolved in isolation (simulation Q from Tab. 1), we show how the projected logarithmic density profile slope, $\gamma_{\mathrm{2D}}$, and the projected enclose mass, $M_{<r_{\mathrm{2D}}}$, evolve with time. We display $\gamma_{\mathrm{2D}}$ and $M_{<r_{\mathrm{2D}}}$ for several radii $r_{\mathrm{2D}}$ as indicated in the legend. The larger circles mark the value at the beginning of the simulation, i.e. for an NFW profile. Moreover, the small white dots are placed equidistant in time, every 0.98 Gyr.
[…]
@@
Figure 11: Enclosed mass within several radii as a function of time. The results are shown for simulations with the standard time step criterion (darker curves) and those employing a minimum time step (lighter curves). The three panels give the enclosed mass within different radii for the same simulations as shown in Fig. 7. Moreover, we show the enclosed mass for the initial NFW halo as a reference (grey).
@@
In addition to the results above, we also show how the enclosed mass evolves for our simulations employing a minimum time step. The results for all three of them are given in Fig. 11. It is visible that for the late phase that we would hardly be able to simulate without imposing a minimum time step, the values for the enclosed masses differ significantly between the different runs. As a result, cosmological simulations employing a minimum for the allowed time step may overestimate the collapse time and underestimate the density at the halo centre. But importantly, they may also overestimate the mass of the halo within the scale radius of the initial NFW profile or larger radii. As a consequence, the ability of SIDM simulations to explain gravitational lensing signals supposedly arising from fairly concentrated objects can be easily overestimated.
@@ 3.8 Velocity and angular dependence
Figure 12: Velocity and angular dependent cross-sections. Following Fig. 17, the enclosed mass within 10 pc (upper panel) and the energy conservation (lower panel) are shown. In detail, a velocity-independent isotropic cross-section (blue), a velocity-independent forward-dominated cross-section (light blue), a velocity-dependent isotropic cross-section (red), and a velocity-dependent forward-dominated cross-section (light red) are displayed. They all share the same resolution of $N=10^{7}$ and a gravitational softening length of $\epsilon=1\,\mathrm{pc}$. More details on the simulations O, P, R, and S can be found in Tab. 1.
So far, we have discussed the simulation results for an isotropic velocity-independent cross-section only. Now, we will focus on the angular and velocity dependence. In contrast to the previous part, we are mainly interested in qualitative differences arising from particle physics. Given that all the simulations that we will discuss here are based on similar numerical schemes, they also have similar numerical properties.
We simulate cross-sections with different velocity and angular dependencies. For the velocity-dependent cross-sections we use
@@
Here, $\sigma_{\mathrm{V}}$ is the viscosity cross-section as given by Eq. (10). The velocity-dependent cross-sections we simulate are given by $\sigma_{\mathrm{0}}/m=6593.89\,\mathrm{cm}^{2}\mathrm{g}^{-1}$ (in terms of the normalised viscosity cross-section, Eq. (10)) and $w=20\,\mathrm{km}\,\mathrm{s}^{-1}$. All values used for our simulations are given in Tab. 1. Moreover, we simulate two different angular dependencies, isotropic scattering and a forward-dominated cross-section, specifically the limit of keeping the momentum transfer constant while the scattering angle approaches zero (e.g. Kahlhoefer et al., 2014; Fischer et al., 2021a).
In Fig. 12, we display the evolution of the isolated halo with different velocity and angular dependencies. The angular dependencies were matched using the viscosity cross-section (Eq. (10)). As we can see, this results in a similar time evolution for the velocity-independent and dependent cases. Moreover, the energy error is similar as well. This demonstrates that the viscosity cross-section works very well for matching different angular dependencies, as previously stated by Yang & Yu (2022); Sabarish et al. (2024). In contrast to the angular dependence, the velocity dependence (Eq. (4)) leads to a qualitatively different evolution of the halo, with a larger maximum core size and a longer collapse time compared to the core expansion phase. This qualitative behaviour arising from the velocity dependence has been discussed in larger detail in Sect. 3.2.3 by Fischer et al. (2024b). We provided further discussion in Appendix C.
@@ 3.9 Velocity Anisotropy
Figure 13: Velocity anisotropy as a function of radius. We show $\beta$ (Eq. (5)) for highest-resolved simulation (simulation Q of Tab. 1) at different times.
As a last part on the evolution of the isolated halo, we study the velocity anisotropy and compute it as,
@@
with the polar velocity dispersion $\sigma^{2}_{\theta}$, the azimuthal velocity dispersion $\sigma^{2}_{\phi}$, and the radial velocity dispersion $\sigma^{2}_{r}$. The results are displayed in Fig. 13. For $t=0$ the velocity distribution is isotropic at all radii ($\beta=0$) but increases at later times at the larger radii beyond the inner density core. Here, radial motion becomes more pronounced. In the context of SIDM this has been previously pointed out by Gurian & May (2025).
@@ 4 Satellite evolution
In this section, we study the evolution of the SIDM halo under the influence of tidal forces arising from a host galaxy. We first describe the changes to our numerical set-up in Sect. 4.1. An illustration of the satellite orbit and the tidal evolution can be found in Appendix E. We study the difference between a satellite halo and an isolated system (Sect. 4.2) as well as the difference arising from the angular and velocity dependence (Sect. 4.3). Moreover, we demonstrate in Sect. 4.4 that the density profile during the collapse phase can be well described by a King model. Overall, we mainly focus on the physics, in contrast to the numerics of the previous section. Given that we have tested various numerical parameters for the isolated case, we now assume that the same choice also works when evolving the halo in an external potential. We note that we have used the peak find algorithm by Fischer et al. (2021b) to determine the position of the satellite. As in the previous section, we compute physical quantities with respect to this position for all of this section.
@@ 4.1 Numerical set-up
To model the tidal forces of the host system acting on the satellite at low computational costs, we describe the gravitational potential of the host analytically. We only focus on the role of the tidal forces and neglect dynamical friction acting on the satellite and making its orbit decay. Moreover, we also do not take the scattering of DM particles of the satellite halo on the host’s DM into account, which can play an important role especially for velocity-independent cross-sections (e.g. Zeng et al., 2022).
The external potential is chosen to mimic the gravitational potential of the Milky Way and consists of six components. The details of our description follow the one employed by Zhang et al. (2025). We describe the DM halo with an NFW profile (Navarro et al., 1996). The corresponding potential is
@@
Here, G denotes the gravitational constant. For our simulations we set the NFW halo parameters to $\rho_{s}=8.54\times 10^{6}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ and $r_{\mathrm{s}}=19.6\,\mathrm{kpc}$. To mimic the gravitational potential of the stellar bulge we employ a Hernquist profile (Hernquist, 1990).
@@
For our set-up we use $M_{\mathrm{H}}=9.23\times 10^{9}\,\mathrm{M_{\odot}}$ and $a_{\mathrm{H}}=1.3\,\mathrm{kpc}$. On top of this, we add four disk components described by an axisymmetric Miyamoto–Nagai profile (Miyamoto & Nagai, 1975).
$\Phi_{\mathrm{disk}}(r)=-\mathrm{G}\,M_{\mathrm{d}}\left[R^{2}+\left(a_{\mathrm{d}}+\sqrt{z^{2}+b^{2}_{\mathrm{d}}}\right)^{2}\right]^{-1/2}$ (8)
@@
The satellite halo is initially placed at a distance of $r=123.6\,\mathrm{kpc}$ from the centre of the external potential. The exact coordinates are $[58.16,40.12,101.41]\,(\mathrm{kpc})$. The position is relative to the origin of the external potential. The initial velocity of the satellite is $v=109\,\mathrm{km}\,\mathrm{s}^{-1}$. The velocity vector reads $[45.94,83.42,53.12]\,(\mathrm{km}\,\mathrm{s}^{-1})$. With this choice we also follow Zhang et al. (2025). Moreover, we illustrate the orbit of the satellite halo in Appendix E.
4.2 Isolated vs. satellite
@@
Figure 14: Evolution of an isolated halo (solid) and a satellite halo (dashed). The simulations are for a velocity-independent cross-section, employ $N=10^{7}$ particles, and use a gravitational softening length of $\epsilon=1\,\mathrm{pc}$. All parameters for the displayed simulations O and W are given in Tab. 1.
@@
Next, we investigate how the evolution of the density and velocity dispersion profiles changes for a satellite halo compared to a system evolved in isolation. In Fig. 14, we show the evolution of the density, one-dimensional velocity dispersion, and Knudsen number, as a function of radius for a halo in isolation (solid lines) and for a satellite system (dashed lines).
@@
By and large, the evolution of the satellite system is similar to the isolated halo. First, the two undergo a core expansion phase and later collapse. However, the velocity dispersion profile (middle panels of Fig. 14) reveals a significant difference. For the satellite halo, the velocity dispersion is strongly affected by tidal forces. Close to the pericentre passage (e.g. at $1.47\,\mathrm{Gyr}$) tidal heating injects energy into the halo and the velocity dispersion increases. Subsequently particles become unbound and leave the system and the velocity dispersion decreases over all relevant radii. This is in contrast to the isolated halo where the velocity dispersion is only affected by the self-interactions. The speed-up of the halo evolution due to tidal stripping has been found by various authors (e.g. Kahlhoefer et al., 2019; Sameie et al., 2020). This makes the energy outflow of the satellite more efficient and speeds up the halo evolution, i.e. reduces the collapse time. Later, in the collapse phase, the velocity dispersion in the centre reaches unprecedentedly high values.
@@
We note, the satellite system reaches a central density of $\rho=3\times 10^{11}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ in the last snapshot at $t=4.6\,\mathrm{Gyr}$. At sufficiently small radii, the density gradient becomes flat, and the Knudsen number reaches a value of about 0.04, i.e. the central region is deep in the short-mean-free-path regime. In contrast, the lowest central density we find at maximum core expansion is about $\rho=6\times 10^{7}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$. It is lower compared to the isolated case due to tidal interactions. Moreover, the tidal interactions also greatly reduce the density in the outer regions of the halo and lead to a large loss in gravitationally bound mass.

### v4

@@ 3.7 Evolution of projected enclosed mass
Figure 9: Time evolution of the enclosed mass for our most highly resolved simulation (simulation Q in Table 1). The relative change of the enclosed mass within various radii is shown as a function of time. The enclosed mass is computed in projection, i.e. it is the mass within a cylinder of radius $r_{\mathrm{2D}}$.
@@
We wanted to get closer to the observations, in particular gravitational lensing, and to measure the enclosed mass in projection. In Fig. 9, the enclosed mass within several radii is shown as a function of time for our most highly resolved simulation ($N=5\times 10^{7}$). For small radii, we find that the enclosed mass first decreases during the core expansion phase and later increases when the gravothermal collapse dominates, as we also found in our three-dimensional analysis. This is in contrast to the bound mass of the system. The self-interactions create a few particles that have velocities exceeding the escape velocity, leading to a decreasing mass of the halo. Even more particles may gain enough kinetic energy to travel to fairly large radii, but still remain bound. In line with this, we find that for the largest radius shown in Fig. 9, the mass decreases slightly. Related to this, the velocity anisotropy at large radii increases, i.e. radial motion becomes more important, as we show in Appendix G.
@@
If one defines a radius $r_{\mathrm{eq}}$ as the one at which the mass inflow and the mass outflow have the same size, i.e. $\dot{M}(<r_{\mathrm{eq}})=0$. Then $r_{\mathrm{eq}}$ decreases with time when the halo is undergoing the gravothermal collapse. This implies that we do not expect the mass within a fixed radius (e.g. 1 kpc) to keep continuously increasing. Instead, it starts to decrease as we can see for a few of the radii that we have chosen for Fig. 9. The smaller the radius, the later the point in time where the enclosed mass starts to decrease. [^1] For a more complicated model, such as the one studied by Patil & Fischer (2025), the decrease in mass in the late stages can be avoided by approaching a stable solution with an increased enclosed mass compared to collisionless DM instead of undergoing a gravothermal catastrophe.
@@
Figure 10: Projected logarithmic density profile slope as a function of projected enclosed mass. For our most highly resolved simulation evolved in isolation (simulation Q from Table 1), we show how the projected logarithmic density profile slope, $\gamma_{\mathrm{2D}}$, and the projected enclose mass, $M_{<r_{\mathrm{2D}}}$, evolve with time. We display $\gamma_{\mathrm{2D}}$ and $M_{<r_{\mathrm{2D}}}$ for several radii $r_{\mathrm{2D}}$ (see inset for legend). The larger circles mark the value at the beginning of the simulation, i.e. for an NFW profile. Moreover, the small white dots are placed equidistant in time, every 0.98 Gyr.
The projected logarithmic density profile slope, $\gamma_{\mathrm{2D}}$, shows a similar behaviour. In Fig. 10, we give $\gamma_{\mathrm{2D}}$ as a function of the enclosed mass $M_{<r_{\mathrm{2D}}}$. The two quantities can be inferred from observations via gravitational lensing analysis. Here we use our most highly resolved simulation, which models the halo in isolation. The figure shows how the system evolves in the $\gamma_{\mathrm{2D}}$$M_{<r_{\mathrm{2D}}}$ plane considering several radii, $r_{\mathrm{2D}}$. It can be seen that not only does the enclosed mass make a turn during the collapse phase, as discussed above, but the density slope also starts to become flatter at a specific point in time. Similarly to the enclosed mass, this turn occurs earlier at larger radii. Interestingly, $\gamma_{\mathrm{2D}}$ turns even earlier to becoming flatter than the projected enclosed mass starts to decrease.
@@
Figure 11: Enclosed mass within several radii as a function of time. The results are shown for simulations with the standard time step criterion (darker curves) and those employing a minimum time step (lighter curves). The three panels give the enclosed mass within different radii for the same simulations, as shown in Fig. 7. Moreover, we show the enclosed mass for the initial NFW halo as a reference (grey).
@@
In addition to the results above, we also show how the enclosed mass evolves for our simulations employing a minimum time step. The results for all three of them are given in Fig. 11. The figure shows that for the late phase that we would hardly be able to simulate without imposing a minimum time step; the values for the enclosed masses differ significantly between the different runs. As a result, cosmological simulations employing a minimum for the allowed time step may overestimate the collapse time and underestimate the density at the halo centre. But importantly, they may also overestimate the mass of the halo within the scale radius of the initial NFW profile or larger radii. As a consequence, the ability of SIDM simulations to explain gravitational lensing signals supposedly arising from fairly concentrated objects can be easily overestimated.
@@ 3.8 Velocity and angular dependence
Figure 12: Velocity and angular dependent cross-sections. Following Fig. 1, the enclosed mass within 10 pc (upper panel) and the energy conservation (lower panel) are shown. In detail, a velocity-independent isotropic cross-section (blue), a velocity-independent forward-dominated cross-section (light blue), a velocity-dependent isotropic cross-section (red), and a velocity-dependent forward-dominated cross-section (light red) are displayed. They all have the same resolution of $N=10^{7}$ and a gravitational softening length of $\epsilon=1\,\mathrm{pc}$. More details on the simulations O, P, R, and S can be found in Table 1.
So far, we have discussed the simulation results for an isotropic velocity-independent cross-section only. Now we focus on the angular and velocity dependence. In contrast to the previous part, we are mainly interested in qualitative differences arising from particle physics. Given that all the simulations we discuss here are based on similar numerical schemes, they also have similar numerical properties.
We simulated cross-sections with different velocity and angular dependences. For the velocity-dependent cross-sections we used
@@
Here, $\sigma_{\mathrm{V}}$ is the viscosity cross-section as given by Eq. (10). The velocity-dependent cross-sections we simulated are given by $\sigma_{\mathrm{0}}/m=6593.89\,\mathrm{cm}^{2}\mathrm{g}^{-1}$ (in terms of the normalised viscosity cross-section, Eq. (10)) and $w=20\,\mathrm{km}\,\mathrm{s}^{-1}$. All values used for our simulations are given in Table 1. Moreover, we simulated two different angular dependences, isotropic scattering and a forward-dominated cross-section, specifically the limit of keeping the momentum transfer constant while the scattering angle approaches zero (e.g. Kahlhoefer et al., 2014; Fischer et al., 2021a).
In Fig. 12, we display the evolution of the isolated halo with different velocity and angular dependences. The angular dependences were matched using the viscosity cross-section (Eq. (10)). As we can see, this results in a similar time evolution for the velocity-independent and dependent cases. Moreover, the energy error is similar as well. This demonstrates that the viscosity cross-section works very well for matching different angular dependences, as previously stated by Yang & Yu (2022); Sabarish et al. (2024). In contrast to the angular dependence, the velocity dependence (Eq. (4)) leads to a qualitatively different evolution of the halo, with a larger maximum core size and a longer collapse time compared to the core expansion phase. This qualitative behaviour arising from the velocity dependence is discussed in greater detail in Fischer et al. (2024b) (Sect. 3.2.3). We provide further discussion in Appendix C.
@@ 3.9 Velocity anisotropy
Figure 13: Velocity anisotropy as a function of radius. We show $\beta$ (Eq. (5)) for our most highly resolved simulation (simulation Q of Table 1) at different times.
As a last part on the evolution of the isolated halo, we present the velocity anisotropy, and compute it as
@@
with the polar velocity dispersion $\sigma^{2}_{\theta}$, the azimuthal velocity dispersion $\sigma^{2}_{\phi}$, and the radial velocity dispersion $\sigma^{2}_{r}$. The results are displayed in Fig. 13. For $t=0$ the velocity distribution is isotropic at all radii ($\beta=0$), but increases at later times at the larger radii beyond the inner density core. Here, the radial motion becomes more pronounced (see also Gurian & May (2025)). We note that at the largest radii shown in Fig. 13, the anisotropy increases. This might be a consequence of the particles moving radially outwards becoming more dominant as the density decreases strongly with radius ($\rho\propto r^{-3}$). In addition, the collision rate is low at those radii, and the self-interactions do not efficiently isotropise the velocity distribution.
@@ 4 Satellite evolution
In this section we discuss the evolution of the SIDM halo under the influence of tidal forces arising from a host galaxy. We first describe the changes to our numerical set-up in Sect. 4.1. An illustration of the satellite orbit and the tidal evolution can be found in Appendix E. We show the difference between a satellite halo and an isolated system (Sect. 4.2) as well as the difference arising from the angular and velocity dependence (Sect. 4.3). Moreover, we demonstrate in Sect. 4.4 that the density profile during the collapse phase can be well described by a King model. Overall, we mainly focus on the physics, in contrast to the numerical aspects of the previous section. Given that we tested various numerical parameters for the isolated case, we now assume that the same choice also works when evolving the halo in an external potential. We note that we used the peak find algorithm by Fischer et al. (2021b) to determine the position of the satellite. As in the previous section, we compute physical quantities with respect to this position for all of this section.
@@ 4.1 Numerical set-up
To model the tidal forces of the host system acting on the satellite at low computational costs, we described the gravitational potential of the host analytically. We focused only on the role of the tidal forces and neglected the dynamical friction that acts on the satellite and causes its orbit to decay. Moreover, we also did not take the scattering of DM particles of the satellite halo on the host’s DM into account, which can play an important role especially for velocity-independent cross-sections (e.g. Zeng et al., 2022).
The external potential is chosen to mimic the gravitational potential of the Milky Way and consists of six components. The details of our description follow that employed by Zhang et al. (2025). We describe the DM halo with an NFW profile (Navarro et al., 1996). The corresponding potential is
@@
Here G denotes the gravitational constant. For our simulations, we set the NFW halo parameters to $\rho_{s}=8.54\times 10^{6}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ and $r_{\mathrm{s}}=19.6\,\mathrm{kpc}$. To mimic the gravitational potential of the stellar bulge, we employed a Hernquist profile (Hernquist, 1990).
@@
For our set-up, we used $M_{\mathrm{H}}=9.23\times 10^{9}\,\mathrm{M_{\odot}}$ and $a_{\mathrm{H}}=1.3\,\mathrm{kpc}$. In addition, we added four disk components described by an axisymmetric Miyamoto–Nagai profile (Miyamoto & Nagai, 1975),
$\Phi_{\mathrm{disk}}(r)=-\mathrm{G}\,M_{\mathrm{d}}\left[R^{2}+\left(a_{\mathrm{d}}+\sqrt{z^{2}+b^{2}_{\mathrm{d}}}\right)^{2}\right]^{-1/2},$ (8)
@@
The satellite halo is initially placed at a distance of $r=123.6\,\mathrm{kpc}$ from the centre of the external potential. The exact coordinates are $58.16,40.12,101.41]\,(\mathrm{kpc})$. The position is relative to the origin of the external potential. The initial velocity of the satellite is $v=109\,\mathrm{km}\,\mathrm{s}^{-1}$. The velocity vector reads $45.94,83.42,53.12]\,(\mathrm{km}\,\mathrm{s}^{-1})$. With this choice, we also follow Zhang et al. (2025). Moreover, we illustrate the orbit of the satellite halo in Appendix E.
4.2 Isolated versus satellite
@@
Figure 14: Evolution of an isolated halo (solid line) and a satellite halo (dashed line). The simulations are for a velocity-independent cross-section, employ $N=10^{7}$ particles, and use a gravitational softening length of $\epsilon=1\,\mathrm{pc}$. All parameters for the displayed simulations O and W are given in Table 1.
@@
Next, we discuss how the evolution of the density and velocity dispersion profiles changes for a satellite halo compared to a system evolved in isolation. In Fig. 14, we show the evolution of the density, the one-dimensional velocity dispersion, and the Knudsen number, as a function of radius for a halo in isolation (solid lines) and for a satellite system (dashed lines).
@@
By and large, the evolution of the satellite system is similar to that of the isolated halo. First, the two undergo a core expansion phase and later collapse. However, the velocity dispersion profile (middle panels of Fig. 14) reveals a significant difference. For the satellite halo, the velocity dispersion is strongly affected by tidal forces. Close to the pericentre passage (e.g. at $1.47\,\mathrm{Gyr}$) tidal heating injects energy into the halo and the velocity dispersion increases. Subsequently, particles become unbound and leave the system, and the velocity dispersion decreases over all relevant radii. This is in contrast to the isolated halo, where the velocity dispersion is only affected by the self-interactions. The speed-up of the halo evolution due to tidal stripping has been found by various authors (e.g. Kahlhoefer et al., 2019; Sameie et al., 2020). This makes the energy outflow of the satellite more efficient and speeds up the halo evolution, i.e. reduces the collapse time. Later, in the collapse phase, the velocity dispersion in the centre reaches unprecedentedly high values.
@@
We note that the satellite system reaches a central density of $\rho=3\times 10^{11}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ in the last snapshot at $t=4.6\,\mathrm{Gyr}$. At sufficiently small radii, the density gradient becomes flat and the Knudsen number reaches a value of about 0.04, i.e. the central region is deep in the short-mean-free-path regime. In contrast, the lowest central density we find at the maximum core expansion is about $\rho=6\times 10^{7}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$. It is lower compared to the isolated case due to tidal interactions. Moreover, the tidal interactions also greatly reduce the density in the outer regions of the halo, which leads to a large loss in gravitationally bound mass.